Home‎ > ‎Reports‎ > ‎crew114‎ > ‎

day04


Crew #114 Log Book for February 28, 2012

Photos of the day





Commander's Check-In Report


Commander's Report

Date and time: February 28, 2012; 19:00

Written by: Balwant Rai, Crew 114, Commander

Today we started with breakfast and briefing the day. Afterwards our Health and safety officer examined our dental health. Today, we started with Rover and geology experiments followed by medical experiments. We had good lunch prepared by Commander and Chief executive officer. This was followed by two EVAs that involved collecting soil samples for analysis of microbiological activity.


During the late afternoon, the cadets worked on getting the rover working and examining how well it maneuvers when being controlled outside of the Hab. Our chefs are currently working on a good cooked meal tonight, since it is a cooking-day.


Tomorrow, we plan to begin doing the medical experiments, especially dental and human factors experiment again on all crew members. The rover engineer, Yannick Wable, is planning on continuing his study of the rover and potentially using the drone if the environment conditions allow. Pierrick and Thomas Quemin will also join Pierrick in examining the feasibility of the rover here on Mars. Wouter Poos will continue his collection of soil samples. Medical data will be collected during two EVAs tomorrow by Balwnt Rai and Jasdeep Kaur. One will be on foot, and one will use the ATVs. Tomorrow will be a cooking day, so we look forward for hot and very tasty meals.

Science Report (Geology)

Date and time: February 28, 2012; 11:00

Written by: Wouter Poos, Crew 114, Crew Geologist

Took surface and superficial microbiological samples at site 1 during EVA 1. No problems encountered. Samples are stored in the refrigerator for further analysis.

Science Report (Geology) 2

Date and time: February 28, 2012 13:30

Written by Wouter Poos, Crew 114, Crew Geologist:

Took subsurface samples (0, 25, 50 cm deep) of brushy basin member at three different sites, one on alluvial plain, one at base of slope and one in canyon. Successful. Samples are stored in Lab for further analysis.

Science Report (Rover)


Date and time: February 28, 2012; 19:00

Written by: Wable Yannick, Crew 114, Rover engineer

Crew Members: Every member of the crew.
Description: The objective of this experiment was to improve the efficiency of rover driving skill and to test new members. Objective has been achieved

Engineering Report

Date and time: February 27, 2012, 16:30
Written by Pierrick DARNEAUX, Crew 114, Crew Engineer.

Generator/Electricity:
Did you have any electricity outage? (Yes/No): NO

Fuel Status:
Diesel (Full/0.75/0.5/0.25/Empty): <0.25
Propane Tank (Full/0.75/0.5/0.25/Empty): N/A
Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.5

Oil:
ATV Oil Quantity (# of Quarts in storage): 20
Car Oil Quantity (# of Quarts in storage): N/A

Water Status:
Outside Potable Water Tank Level (inches from bottom): 11
Trailer Potable Water Tank Level (inches from bottom): 14
Hab Potable Water Tank Level (inches from bottom): 11
Potable Water Meter Reading: 0051570.4
Into GreenHab Meter Reading: N/A
Notes/Comments:

GreenHab:
Condition of plants in Tank 1: N/A
Condition of plants in Tank 2: N/A
Condition of plants in Tank 3: N/A

Observatory:
Telescope Used? (Yes/No): NO
Observatory Notes/Comments: N/A

Transportation:
Hab Car (NOT RENTAL)
Used (yes/no): N/A
Oil Checked (yes/no): YES: N/A
Tire Status: N/A
Odometer at end of day: N/A
Notes/Comments on Hab Car: No car

ATV 1 (Opportunity):
Used: YES
Oil Checked: YES: Full
Fuel Consumed: 0.5gal
Tire Status: GOOD
Today's Run Time: 30min
Season's Run Time: 88:25
Comments:-

ATV 2 (Spirit):
Used: NO
Oil Checked: YES: Full
Fuel Consumed: 0
Tire Status: GOOD
Today's Run Time: 0
Season's Run Time: 82:55
Comments:-

ATV 3 (Viking I):
Used: YES
Oil Checked: YES: Full
Fuel Consumed: 0.5gal
Tire Status: GOOD
Today's Run Time: 0.30min
Season's Run Time: 86:25
Comments:-

Heating and Ventilation:
Thermostat setting upstairs (Farenheit):69°F
Actual temperature upstairs (Farenheit):74°F

Thermostat setting downstairs (Farenheit):65°F
Actual temperature downstairs (Farenheit):64°F

Computers/Networking Infrastructure:
Notes/comments on internet/computers?:

General Engineering Notes/Comments: N/A


EVA 5 Report

Date and time: February 28, 2012 17:27
Written by: Pierrick Darneaux, Crew 114, Engineer
Crew Members: Geologist (Woulter Poos), health and safety officer (Jasdeep Kaur), Rover engineer (Yannick Wable) and engineer (Perrick Darneaux)
Site Location : UTM UPS (Location System) Classified
Session: 10:00-11:10
Objective: Sample collecting mission
Results: Successful
Lesson learned: How to take a rock sample

Chef's Report

Date and Time: February 28, 2012; 19:00

Written by: Quemin Thomas, Crew 114,Executive officer

Chefs: commander and the executive officer


1. Today was

a cooking day

2. 1b) Was today a special day celebrated at the main meal with special food or activity? If so, what was special about today and what food(s) or activities were prepared to mark the celebration?

No

3. Description:

Potato Flakes / eggs / boiled water / bacon (a whole pack)/ spices



4. “Special Potato Flakes and spatio omelet”
a. 30 oz of potato Flakes
Add 6 cups of boiled water
Add salt + chili powder (just a little) + “herbes de provence”
then take care of it during 15 minutes

b. In a frying pan put some oil and add 3 spoons of onions with salt.
Melt 6 tablespoons of powdered eggs and add 9 spoons of water (in a bowl)
Cut the bacon in little pieces

c. When the onions are ready add the eggs and bacon in the frying. Cook it exactly 7min and 32 seconds.

“Enjoy your meal!”

5. List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

Nothing

6. Lessons learned : you may multiply by 3 the quantities for the omelet making (not for bacon added)